OpenStudy (anonymous):

Find the value of k so that the graph of y=(x+5)^2+k passes through the point (0,−12).

OpenStudy (mayankdevnani):

so x=0 and y=-12

OpenStudy (anonymous):

Just substitute in (0,-12) into the equation... and then solve for k y = (x + 5)^2 + k -12 = (0 + 5)^2 + k -12 = 5^2 + k Solve for k...

OpenStudy (mayankdevnani):

-12=(0+5)^2+k -12=25+k k=-25-12=- 37

OpenStudy (mayankdevnani):

@1193ped ok

OpenStudy (anonymous):

thanks man